# SAMSA: Efficient Transformer for Many Data Modalities

### v1

@@ 5.1. Experimental Design
SAMSA-hard (ours) 0.2458 ± 0.0013 0.7221 ± 0.0042
SAMSA-soft (ours) 0.2519 ± 0.0011 0.7171 ± 0.0051
We hypothesize that SAMSA’s superior performance can be attributed to the regularization effect introduced by the sampling operator. During the early stages of training, the sampling mechanism tends to select tokens almost randomly, which is akin to the DropAttention approach (Zehui et al., 2019). This random selection acts as a form of regularization, preventing the model from overfitting by ensuring that attention is not overly focused on any particular subset of tokens. As training progresses and the sampler becomes more adept at identifying important tokens, the attention mechanism becomes more targeted, gradually diminishing the regularization effect. This dynamic behavior aligns with previously observed phenomena, where early-stage regularization accelerates the model’s convergence rate, leading to better overall performance (Liu et al., 2023b).
Regarding the trade-off between performance and efficiency, the data in Table 1 reveals a significant insight: SAMSA can achieve superior results with a sampling rate of just 256 tokens per head, outperforming both full self-attention and configurations with higher sampling rates. This finding strongly suggests that not all tokens contribute equally to the attention mechanism’s effectiveness. Instead, there exists an optimal or ”Goldilocks-zone” sampling rate where the regularization effect of sampling maximizes performance while simultaneously reducing inference time. This balance between efficiency and effectiveness is a key advantage of SAMSA, as it highlights the potential to streamline attention computation without sacrificing accuracy.
Table 4. Experimental results on ModelNet40 dataset (Wu et al., 2015) and part segmentation results on ShapeNetPart dataset (Chang et al., 2015) against point-cloud specialized baselines.
@@
SAMSA-hard (Ours) 88.9 91.3 83.8 85.8
SAMSA-soft (Ours) 88.7 91.7 - -
5.3. Sampling vs. Specialized Models
A significant advantage of SAMSA is its applicability across a wide range of data modalities, making it a versatile tool in diverse fields. This versatility is evident in the performance comparisons shown in Table 2, Table 3, and Table 4, where SAMSA is evaluated against a variety of specialized methods in graph processing (using the Long Range Graph Benchmark) and point cloud analysis (using ModelNet40 and ShapeNetPart). Despite SAMSA being designed without any inherent inductive biases tailored to specific data types, it demonstrates a performance that is not only comparable to these specialized methods but also surpasses many graph-based models.
In the context of graph processing, SAMSA produced competitive results on both the Peptides Func and Struct datasets. Notably, it achieved an average precision (AP) of $72.21\%$, outperforming all other graph methods evaluated in these benchmarks. This highlights SAMSA’s ability to effectively capture and utilize the important features within graph data, even when competing against models specifically designed for such tasks.
For point cloud analysis, while SAMSA did not reach the same level of performance as in graph processing, primarily due to the absence of local information aggregation mechanisms which are crucial in this domain, it still delivered results that were on par with many models specifically optimized for point clouds. This demonstrates SAMSA’s robustness and adaptability, as it remains competitive even in scenarios where it faces inherent challenges. Overall, SAMSA’s ability to perform well across different data modalities underscores its potential as a foundational model, capable of replacing or complementing more specialized approaches in various applications.
6. Conclusion
In conclusion, the SAMSA (SAMpling-based Self-Attention) mechanism addresses the limitations of traditional transformers by reducing the quadratic complexity typically associated with self-attention. Unlike other efficient transformers that are often data-modality-specific, SAMSA provides a versatile, context-aware linear complexity solution that can be applied across various data modalities. By employing a differentiable sampling method to focus on the most important tokens, SAMSA achieves competitive or even state-of-the-art results on multiple benchmarks. Additionally, SAMSA’s sparse formulation during inference minimizes computational costs, offering faster inference times with little in performance, and in some cases, even outperforming full self-attention transformers.
@@ 6. Conclusion
Despite the promising results achieved with SAMSA, there are certain limitations on the biased gradients of hard sampling formulation, which makes our model hard to converge in more difficult problems (Pathfinder). The biased gradient issue in hard sampling methods is mostly from the misestimation of gradient of the non-chosen path. We believe that future work on more precise gradient estimation would yield inference-fast model with much better accuracy.
@@ Appendix A Datasets
The Long Range Arena (LRA) (Tay et al., 2021) is a composition benchmark of 5 tasks: ListOps (Nangia and Bowman, 2018), ImDB review (Maas et al., 2011), ACL Anthology Network (Radev et al., 2009), Grayscaled CIFAR-10 (Krizhevsky and Hinton, 2009), and Pathfinder (Linsley et al., 2018). These five tasks are all classification and they feature very long sequences: ListOps (2,048 tokens), ImDB review (1,024 tokens), ACL Anthology Network (4,096 tokens), Grayscaled CIFAR-10 (1,024 tokens), and Pathfinder (1,024 tokens). All five tasks involves tackling long-ranged data structures and manifold categorizations, challenging networks’s generalization powers as well as memory and time efficiencies.
@@ Appendix B Training Strategies and Data Preprocessing
Here is the list of hyperparameters:
  • Optimizer: AdamW (Loshchilov and Hutter, 2019)
@@
  • Learning rate warm-up: Untuned Linear Warmup (Ma and Yarats, 2021)
@@
Table 5. Training details.
Task Dataset Batch Size No. of Epochs
Point Cloud ModelNet40 32 600
ShapeNetPart 32 600
Sequence ListOPS 32 50
Text 32 50
Retrieval 32 50
Image 32 50
Pathfinder 32 50
Graph Peptides-struct 32 240
Peptides-func 32 160
@@
  • ModelNet40: Following previous work (Charles et al., 2017), we use farthest point sampling to sample 1024 initial points.
@@
All of the training processes utilize Cross Entropy Loss, without label smoothing normalization.

### v2

@@ 5.2 Performance-Efficiency Trade-offs
24-layer hard-SAMSA (ours) 0.2458 ± 0.0013 0.7221 ± 0.0042
24-layer soft-SAMSA (ours) 0.2519 ± 0.0011 0.7171 ± 0.0051
In Table 3, our model has outperformed most methods in the two tasks we tested. This shows the efficacy of our newly proposed Graph-PE. Note that, our method’s learnable positional encoding is computed cheaply, and how cheap the computation is best represented in Table 4. The table shows that our entire model computes everything even before other methods finish their positional encoding computation. It also shows the efficacy of the SAMSA model as a very deep network, which is also different from traditional graph learning methods which cannot extend beyond 8 layers due to over-smoothing.
Table 4: Runtime of Graph Positional Encoding on one epoch on A100 compared to our 24-layer SAMSA model. The runtime of LapPE and RWPE are taken from (Dwivedi et al., 2022).
Method LapPE RWPE SAMSA (whole model)
Peptides Struct 73s 53s 35s
Table 5: Inference time (in seconds) of the networks for ModelNet40 classification test split in 1 A100 and 8 CPUs with a batch size of 32. The results of PointNet and DGCNN are taken from (Kaba et al., 2023).
Method PointNet DGCNN hard-SAMSA (ours) soft-SAMSA (ours)
Inference Time 18s 23s <1s 1s
Table 6: Experimental results on ModelNet40 dataset (Wu et al., 2015) and part segmentation results on ShapeNetPart dataset (Chang et al., 2015) against point-cloud specialized baselines.
@@
hard-SAMSA (Ours) 88.9 91.3 83.8 85.8
soft-SAMSA (Ours) 88.7 91.7 - -
Point Cloud.
Performance-wise, Table 6 shows that SAMSA has competitive results against other point cloud models. This is mainly because our model does not include any data-modality-specific inductive bias (like the usage of kNN for locality heuristic); therefore, it is susceptible to overfitting in scenarios where data is not abundant. Our method is also much faster than others, up to 18$\times$ faster than PointNet (which is the most small, simple, efficient point cloud model), as shown in Table 5.
6 Conclusion
SAMSA (SAMpling-based Self-Attention) mechanism addresses the limitations of traditional transformers by reducing the quadratic complexity typically associated with self-attention. Unlike other efficient transformers that are often data-modality-specific, SAMSA provides a versatile, context-aware linear complexity solution that can be applied across various data modalities. By employing a differentiable sampling method to focus on the most important tokens, SAMSA achieves competitive or even state-of-the-art results on multiple benchmarks. Additionally, SAMSA’s sparse formulation during inference minimizes computational costs, offering faster inference times with little in performance, and in some cases, even outperforming full self-attention transformers.
@@ 6 Conclusion
Despite the promising results achieved with SAMSA, there are certain limitations on the misaligned gradients of hard sampling formulation, which makes our model hard to converge in more difficult optimization problems (Pathfinder). This misaligned gradient issue in hard sampling methods is mostly from the misestimation of gradients of the non-chosen path. We believe that future work should focus on more precise gradient estimation would result in an inference-fast model with much better accuracy.
@@ Appendix A Datasets
The Long Range Arena (LRA) (Tay et al., 2021) is a composition benchmark of 5 tasks: ListOps (Nangia & Bowman, 2018), ImDB review (Maas et al., 2011), ACL Anthology Network (Radev et al., 2009), Grayscaled CIFAR-10 (Krizhevsky & Hinton, 2009), and Pathfinder (Linsley et al., 2018). These five tasks are all classification and they feature very long sequences: ListOps (2,048 tokens), ImDB review (1,024 tokens), ACL Anthology Network (4,096 tokens), Grayscaled CIFAR-10 (1,024 tokens), and Pathfinder (1,024 tokens). All five tasks involves tackling long-ranged data structures and manifold categorizations, challenging networks’s generalization powers as well as memory and time efficiencies.
@@ Appendix B Training Strategies and Data Preprocessing
Here is the list of the common hyperparameters:
  • Optimizer: AdamW (Loshchilov & Hutter, 2019)
  • Batch size: 32
@@
  • Learning rate warm-up: Linear Warmup-2000 steps
@@
Table 7: Training hyperparameter details.
Task Dataset Learning Rate Gradient Clipping Weight Decay No. of Epochs
Point Cloud ModelNet40 0.002 2.0 0.1 200/600
ShapeNetPart 0.002 2.0 0.1 600
Sequence ListOPS 0.001 2.0 0.01 50
Text 0.001 2.0 0.01 50
Retrieval 0.001 2.0 0.01 50
Image 0.001 2.0 0.01 50
Pathfinder 0.001 2.0 0.01 50
Graph Peptides-struct 0.001 2.0 0.01 240
Peptides-func 0.0004 2.0 0.01 160
@@
  • ModelNet40: Following previous work (Charles et al., 2017), we use farthest point sampling to sample 1024 initial points with normals.
@@
[…]
Appendix C Architectural Specification
Table 8 shows our model’s architectural hyperparameters, for reproducibility purposes.
Table 8: Architectural details.
Dataset $\mathbf{n_{\textbf{depth}}}$ $\mathbf{d_{\textbf{model}}}$ $\mathbf{d_{\textbf{FFN}}}$ $\mathbf{n_{\textbf{heads}}}$ $\mathbf{p_{\textbf{dropout}}}$ $\mathbf{p_{\textbf{droppath}}}$
ModelNet40 6 256 768 8 0.1 0.1
ShapeNetPart 8 256 768 8 0.1 0.1
LRA-ListOPS 6 256 768 8 0.1 0.0
LRA-Pathfinder 6 128 512 8 0.3 0.0
LRA-Text 4 128 512 8 0.3 0.0
LRA-Image 8 128 512 8 0.3 0.0
LRA-Retrieval 6 128 512 8 0.3 0.0
LRGB-Peptides-Func 24 128 512 8 0.3 0.2
LRGB-Peptides-Struct 24 128 512 8 0.3 0.2
LRA Tasks use Sinusoid-1D positional encodings.
Appendix D Miscellaneous
ST-Gumbel-Sigmoid’s Formula:
$\operatorname{GumbelSigmoid}(\pi)$ $\triangleq\frac{1}{1+\exp\left((\pi_{j}+g_{j})/\tau\right)},$
$\operatorname{ST-GumbelSigmoid}(\pi)$ $\triangleq 1\begin{cases}1,&\text{if }\pi\leq 0,\\ 0,&\text{otherwise},\end{cases}$
$\partial{\operatorname{ST-GumbelSigmoid}}/\partial{\pi}$ $\triangleq\partial{\operatorname{GumbelSigmoid}}/\partial{\pi},$
where $g_{i}$ are independent and identically samples from Gumbel(0,1) distribution, $\pi_{i}$ are the unnormalized logits, and $\tau$ is the temperature parameter that controls the smoothness of the distribution.
Appendix E Learning Curves of Our Models
Figure 2,  3,  4 show the learning curve of SAMSA in various settings, providing for future understanding of how our model learned. All of the lines are plotted with EMA-5 smoothing.
Figure 2: Learning Curves of SAMSA models in Sequence Tasks
Figure 3: Learning Curves of SAMSA models in Graph Tasks
Figure 4: Learning Curves of SAMSA models in Point Cloud Tasks